OpenStudy (loser66):

calculus 5, help please Check directly near which points, we can solve the equation F(x,y) = y^2+y+3x+1 =0

OpenStudy (loser66):

@xapproachesinfinity

OpenStudy (loser66):

I got \(y = -\dfrac{1}{2}\pm\dfrac{\sqrt3}{2}\sqrt{-1-4x}\)

OpenStudy (xapproachesinfinity):

hmm how?

OpenStudy (loser66):

just quadratic equation for y

OpenStudy (xapproachesinfinity):

oh i see. so you are saying any two points satisfying that solution is good!?

OpenStudy (loser66):

I think the solution is the whole parabola. However, if \(x= \dfrac{-1}{4}\), y = -1/2, then \(F(x,y)\neq 0\)

OpenStudy (loser66):

My question: 1) How to argue that the whole parabola is the solution but that point? 2) Are there any other points? How to check?

OpenStudy (xapproachesinfinity):

hmm what if your method was wrong lol

OpenStudy (xapproachesinfinity):

how about if you differentiate

OpenStudy (loser66):

Read the question, kid. It says :"check directly"

OpenStudy (loser66):

If you want to take differentiate, hold on, next problem, we will have some to play with. hehehehe

OpenStudy (xapproachesinfinity):

hmm well depends what is the meaning of directly here

OpenStudy (xapproachesinfinity):

may be they want us to do trial and error hahaha

OpenStudy (loser66):

ok, let check \(F(x,f(x) )= ( -\dfrac{1}{2}\pm\dfrac{\sqrt3}{2}\sqrt{-1-4x})^2+(-\dfrac{1}{2}\pm\dfrac{\sqrt3}{2}\sqrt{-1-4x})+3x+1\)

OpenStudy (loser66):

is it =0 for all x?

OpenStudy (xapproachesinfinity):

hmm let's see

OpenStudy (xapproachesinfinity):

well apparently not for any x take x=0

OpenStudy (loser66):

it is =0, kid

OpenStudy (xapproachesinfinity):

did you check it

OpenStudy (loser66):

yup

OpenStudy (xapproachesinfinity):

but you are operating in complex set

OpenStudy (loser66):

Let check + one, \( ( -\dfrac{1}{2}+\dfrac{\sqrt3}{2}\sqrt{-1-4x})^2+(-\dfrac{1}{2}+\dfrac{\sqrt3}{2}\sqrt{-1-4x})+3x+1\)

OpenStudy (loser66):

I got 0

OpenStudy (xapproachesinfinity):

here is something i did let x=0 \[f(0,y)=y^2+y=0 \Longrightarrow y=-1 \] disregareded y=0 \[f(x,0)=3x+1=0 \Longrightarrow x=\frac{-1}{3}\] the point (-1/3, -1) works fine but don't how to continue and argue this

OpenStudy (loser66):

hey, if x =0, f(0,y) = y^2+y+1

OpenStudy (xapproachesinfinity):

eh blindness my friend haha it is still the same though y=-1 lol

OpenStudy (xapproachesinfinity):

oh no lol

OpenStudy (xapproachesinfinity):

darn it thought it worked lol

OpenStudy (loser66):

hahaha... kid, some girl took of your soul?? you are here but your heart and your mind follow her. hahaha..

OpenStudy (xapproachesinfinity):

hahaha, kind of lol

OpenStudy (xapproachesinfinity):

eh darn, i give up. why don't use the derivatives, though i don't know how will that be useful

OpenStudy (loser66):

hehehe, thanks any way, kid, want to have something else?

OpenStudy (xapproachesinfinity):

how did you get 0 for x=0 in your equation

OpenStudy (xapproachesinfinity):

i didn't get zero

OpenStudy (xapproachesinfinity):

your equatoin fails for x=0, checked didn't get 0

OpenStudy (loser66):

hey, kid, redo, it is =0

OpenStudy (loser66):

\(( -\dfrac{1}{2}+\dfrac{\sqrt3}{2}\sqrt{-1})^2+(-\dfrac{1}{2}+\dfrac{\sqrt3}{2}\sqrt{-1})+3x+1\)

OpenStudy (xapproachesinfinity):

i'am sure it is not 0 lol

OpenStudy (loser66):

ok, watch what the old man do

OpenStudy (loser66):

\((-\dfrac{1}{2})^2+2(\dfrac{\sqrt3}{2}\sqrt{-1})(-\dfrac{1}{2})+(\dfrac{\sqrt3}{2}\sqrt{-1})^2 -\dfrac{1}{2}+\dfrac{\sqrt3}{2}\sqrt{-1}+1\)

OpenStudy (xapproachesinfinity):

i did that!

OpenStudy (xapproachesinfinity):

old you are just wasting time, it is not zero haha

OpenStudy (xapproachesinfinity):

you missed something there i guess

OpenStudy (xapproachesinfinity):

old man*

OpenStudy (xapproachesinfinity):

I'M CHEATING WITH GOOGLE

OpenStudy (xapproachesinfinity):

i don't even want to look at that page hahaha i will get confused lol things i didn't do yet

OpenStudy (loser66):

http://www.wolframalpha.com/input/?i=%28+- \dfrac{1}{2}%2B\dfrac{\sqrt3}{2}\sqrt{-1}%29^2%2B%28-\dfrac{1}{2}%2B\dfrac{\sqrt3}{2}\sqrt{-1}%29%2B1

OpenStudy (xapproachesinfinity):

hehe nothing appeared

OpenStudy (loser66):

you can copy and paste on Wolfram, it says 0

OpenStudy (loser66):

http://www.wolframalpha.com/input/?i=%28+- \dfrac{1}{2}%2B\dfrac{\sqrt3}{2}\sqrt{-1}%29^2%2B%28-\dfrac{1}{2}%2B\dfrac{\sqrt3}{2}\sqrt{-1}%29%2B1

OpenStudy (loser66):

ha, you can do it by yourself

OpenStudy (xapproachesinfinity):

yes looks like i did square something lol

OpenStudy (xapproachesinfinity):

it is zero

OpenStudy (loser66):

ha!! this stubborn kid!!! hehehe

OpenStudy (xapproachesinfinity):

didn't*

OpenStudy (xapproachesinfinity):

hey old man my mind is not working lol

OpenStudy (xapproachesinfinity):

gotta go study for programming lol

OpenStudy (loser66):

ok, break, I need rest also. thanks for being here

OpenStudy (xapproachesinfinity):

np! good luck with this :)